In May there are 6,000 people classified as unemployed and the size of the labor force...

  1. In May there are 6,000 people classified as unemployed and the size of the labor force is 80,000. Between May and June two events take place: i) 1000 graduates join the labor force and start looking for work, ii) 1000 of the unemployed people get “discouraged” and stop looking for work.

A. Calculate the unemployment rate in May.

B. Assuming that the 1000 graduates have not found jobs by June, what is the unemployment rate in June?

C. Assuming that the 1000 graduates do find jobs by June, what is the unemployment rate in June?
Answer #1

a)

Unemployment rate = Unemployed / Labor Force.

= 6000 /80 000

= 7.5

b)

Discouraged workers are not part of labor force. but another 1000 are added to labor force, so there will not be any change to unemployment ratio:

Unemployment rate = (6000 -1000 +1000) / 80 000 - 1000 - 1000

= 7.5 %

c)

Unemployment rate = 5000/80000

= 6.25 %